Solve following systems by addition
2x+3y=1
5x+3y=16

2x+3y=1
-5x-3y=-16
-3x=-15
x=-5
Double checking to see if these are right?

Answer by elima(1433)

The answer is +5, negative divided by a negative gives you a positive. Your doing everything else correct. Good Job. Just watch your signs.
:)
2x+3y=1
-5x-3y=-16
-3x=-15
x=5